Assume the logarithmic transformation f a utility function, for the consumption of two commodities is given by
ln U = ln4 + 0.5ln X + 0.25lnY
(a) if the price of X is GHS2.50 and that of Y is GHS4.00, calculate the optimal combination for an income of GHS50.00.
1
Expert's answer
2021-07-26T16:39:02-0400

ln U = ln4 + 0.5ln X + 0.25lnY

Solving this equation using logarithms

U = 4X0.5Y0.25

If X is GHS2.50 and Y is GHS4.00 thne

U= 4x2.500.5x 4.00o.25

U=8.9

Solving for Y from the logarithmic transformation of a utility function

ln U = ln4 + 0.5ln X + 0.25lnY

Y= U4 / 256X2

But Y=50 and U = 8.9

50 = 8.94 / 256 X2

X2=(50 X 256)/8.94

X2= 2

X= + 1.4 OR -1.4